OpenStudy (anonymous):

31z - 22 = -11z + 188 Z = ? @Shay17

OpenStudy (anonymous):

31z-22=-11z+188 add 22 31z=-11z+210 add 11z 42z=210 divide 42 z=5

OpenStudy (anonymous):

Medal ? @keiyuchan